OpenStudy (anonymous):

MEDALS! Solve the following system of equations. 4x - 3y + z = -9 x + 2y + 2z = 5 x - 2y + 3z = -2

OpenStudy (anonymous):

do you have an idea where to start?

OpenStudy (texaschic101):

4x - 3y + z = -9 x + 2y + 2z = 5 -->(-4)x + 2y + 2z = 5 --------------- 4x - 3y +z = -9 -4x - 8y - 8z = -20 (result of multiplying by -4) ---------------add 0 - 11y - 7z = -29 -11y - 7z = -29 x + 2y + 2z = 5 -->(-1)x + 2y + 2z = 5 x - 2y + 3z = -2 ---------------- -x - 2y - 2z = -5 (result of multiplying by -1) x - 2y + 3z = -2 ----------------add 0 - 4y + z = -7 -4y + z = -7 -11y - 7z = -29 -4y + z = -7 -->(7)-4y + z = -7 ----------------- -11y - 7z = -29 -28y + 7z = - 49 (result of multiplying by 7) -----------------add -39y + 0 = - 78 -39y = -78 y = -78/-39 y = 2 now sub 2 in for y -4y + z = -7 -4(2) + z = -7 -8 + z = -7 z = 8 - 7 z = 1 now sub 2 in for y and 1 in for z in one of the original equations x + 2y + 2z = 5 x + 2(2) + 2(1) = 5 x + 4 + 2 = 5 x + 6 = 5 x = 5 - 6 x = -1 now lets check it....sub in known values 4x - 3y + z = -9 4(-1) - 3(2) + 1 = -9 -4 - 6 + 1 = -9 -10 + 1 = -9 -9 = -9 (correct) x = -1, y = 2, z = 1 or (-1,2,1) If you have any questions, please feel free to ask :)